OpenStudy (anonymous):

What do both of these functions have in common? functions in comments* A. They have the same horizontal translation B. They have the same vertical shift C. They have the same end behavior D. They have the same vertical stretch

OpenStudy (anonymous):

\[f(x)=\frac{ 1 }{ 2 }\left| 3x-7 \right|-2\] and \[g(x)=1.5\ln (x-2)+7\]

OpenStudy (anonymous):

@phi can you help when you can please